### v2

@@ 2.6 Software and Libraries
This study was conducted with several software tools and libraries for data processing, analysis, and model development. QGIS 3.36.0 was employed for examining satellite imagery and utilizing its Python API to calculate vegetation indices, perform spatial joins, and extract tabular data from the spatial layers. Visual Studio Code version 1.89.1 was utilised as the programming IDE. Python 3.9.18 and a list of its libraries were utilisesed, including the following: matplotlib 3.7.2, seaborn 0.12.2, numpy 1.24.3, pandas 2.0.3, scikit-learn 1.3.0 and scipy 1.11.1.
@@ 3 Results
The hyperparameters and overall accuracies of the final models trained on the entire dataset are reported in Table 5 in Appendix B now that an appropriate estimate of performance has been determined. The outcomes largely align with the previously observed accuracies, which is expected if the test set utilsied was representative of the sample distribution. In this final evaluation, there were no ties for best performance across varieties. SVM-RBF maintained its position as the top-performing model for most varieties and QDA continued to outperform other models for SRA14.
The feature importance analysis for diagnosing RSD revealed some variation between the GB and RF classifiers, however, both generally agreed upon several key findings. Notably, When sugarcane variety was not considered, both classifiers identified DWSI-6 as the most influential feature, and identified DWSI-7, DWSI-2, and spectral bands 2, 5, 6, and 11 among the top 10 most important features. Analysing feature importance for RSD diagnosis with separate models for each sugarcane variety indicated variations between both varieties and classifiers. No single feature appeared in the top 10 importance rankings across all varieties for both classifiers. However, across a majority of instances DWSI-6, DWSI-3, Band 2, 5, 6, 7 and 11 were the most frequently occurring features in the top 10 importance. A complete description of feature importance can be seen in Figure 2 in Appendix C.
@@ 4 Discussion
This study demonstrates that ML algorithms can effectively classify RSD across several varieties with freely available satellite-based spectroscopy data. Consistent with previous studies, RF, SVM, and GB variants demonstrated strong performance (Narmilan et al., 2022; Simões and Rios do Amaral, 2023; Ong et al., 2023).
@@
Despite the high classification accuracies achieved by the models, there were several limitations of the study. Positive class precision was consistently lower than negative class precision for models classifying RSD in Q200 and Q208, indicating a higher rate of false positives across the algorithms. This may have arisen from the assumption that RSD infection in a block was uniformly distributed by labeling all pixels within a block with the overall disease status of that block. This could result in regions with little or no RSD being mislabeled as ’RSD Positive’, thereby affecting model training. Future studies should avoid this assumption where possible and implement a higher resolution sampling method. Interestingly, this observation coincides with the only two varieties identified by Sugar Research Australia as partially resistant to RSD (Sugar Research Australia, 2021, 2024). Partial resistance might reduce the spread of RSD sufficiently to create patches of non-infected areas within otherwise infected paddocks. However, to verify this hypothesis, further research with higher resolution sampling methods would be necessary. Additionally, this study focused on developing models and a prototype system for RSD detection within the Herbert region in February of 2022. Future work should explore the influence of other predictors that could not be accounted for in this study to ensure its reliability and broader applicability in other regions or at other points in time. This includes an investigation into the impacts of temperature, humidity, sunlight duration, flowering, other varieties, simultaneous disease infections and precipitation. RSD detection should be trialed at different stages of the sugarcane growth with time series data to improve its applicability to the industry.
@@ 5 Conclusion
In this study, we demonstrated that RSD can be classified for several varieties with ML algorithms from freely available satellite-based spectroscopy data. The best-performing machine learning algorithm was Support Vector Machine with a Radial Basis Kernel achieving between 85.64% and 96.55% accuracy depending on the variety. The performance of all machine learning algorithms was found to be significantly better than the null distribution for all varieties except SRA14. Inline with current literature, the inclusion of sugarcane variety information and vegetation indices improved classification rates of disease. Additionally, several varieties were found to be more challenging to classify for RSD compared to others. These promising initial results, coupled with the efficiency of classifying 72 blocks within minutes as opposed to months, indicates the potential benefits of implementing a large-scale health monitoring system using satellites and machine learning. Given that RSD is one of the most impactful diseases in sugarcane cultivation, this research significantly advances sugarcane disease management.

### v3

@@ 2.6 Software, Libraries & Experimental Environment
This study was conducted with several software tools and libraries for data processing, analysis, and model development. QGIS 3.36.0 was employed for examining satellite imagery and utilizing its Python API to calculate vegetation indices, perform spatial joins, and extract tabular data from the spatial layers. Visual Studio Code version 1.89.1 was utilised as the programming IDE. Python 3.9.18 and a list of its libraries were utilised, including the following: matplotlib 3.7.2, seaborn 0.12.2, numpy 1.24.3, pandas 2.0.3, scikit-learn 1.3.0 and scipy 1.11.1.
All experiments were conducted on a system running Microsoft Windows 11 Home. The hardware configuration included an Intel(R) Core (TM) i7-8700 CPU operating at 3.20 GHz with 6 cores and 12 logical processors, supported by 16 GB of installed physical memory.
@@ 3 Results
The hyperparameters and overall accuracies of the final models trained on the entire dataset are reported in Table 6 in Appendix B now that an appropriate estimate of performance has been determined. The outcomes largely align with the previously observed accuracies, which is expected if the test set utilised was representative of the sample distribution. In this final evaluation, there were no ties for best performance across varieties. SVM-RBF maintained its position as the top-performing model for most varieties and QDA continued to outperform other models for SRA14.
The feature importance for diagnosing RSD revealed some variation between the GB and RF classifiers, however, both generally agreed upon several key findings. Notably, when sugarcane variety was not considered, both classifiers identified DWSI-6 as the most influential feature, and identified DWSI-7, DWSI-2, and spectral bands 2, 5, 6, and 11 among the top 10 most important features. Analysing feature importance for RSD diagnosis with separate models for each sugarcane variety indicated variations between both varieties and classifiers. No single feature appeared in the top 10 importance rankings across all varieties for both classifiers. However, across a majority of instances DWSI-6, DWSI-3, Band 2, 5, 6, 7 and 11 were the most frequently occurring features in the top 10 importance. A complete description of feature importance can be seen in Figure 7 in Appendix C. The t-test to identify statistically significant vegetation indices generally supports similar findings. A summary of these statistical significances is available in Appendix D.
@@ 4 Discussion
This study demonstrates that ML algorithms can effectively classify RSD across several varieties with freely available satellite-based multispectral data. Consistent with previous studies, RF, SVM, and GB variants demonstrated strong performance (Narmilan et al., 2022; Simões and Rios do Amaral, 2023; Ong et al., 2023). Notably, the SVM with RBF kernel slightly outperformed both RF and GB in this context. This may be attributed to the SVM’s ability to effectively model complex non-linear decision boundaries in high-dimensional settings with limited training samples. Under these conditions tree-based methods may be less effective at capturing complex relationships (Murphy, 2022). Additionally, the underlying structure of the disease-related spectral data may be more effectively separated by a smooth, non-linear decision boundary, potentially contributing to the stronger performance observed with SVM using a RBF kernel. While previous sugarcane disease detection studies have primarily focused on drone-based spectroscopy (Narmilan et al., 2022; Simões and Rios do Amaral, 2023; Moriya et al., 2017), handheld spectroscopy (Grisham et al., 2010; Ong et al., 2023) or hyperspectral satellite data (Apan et al., 2004, 2003), our findings extend this knowledge by demonstrating comparable, if not superior, effectiveness of classifying a sugarcane disease with multispectral satellite data. For example, Simões and Rios do Amaral (2023) reported classification accuracies between 86% and 90% for orange and brown rust with multispectral drone data. In comparison, our study achieved accuracies ranging from 83% and 97% accuracy for RF and SVM, depending on the variety, with freely available multispectral data. This demonstrates that our approach, utilising cost-effective large-scale satellite data, can achieve comparable or superior performance. Moreover, our study outperformed the only previous study detecting asymptomatic sugarcane disease (Grisham et al., 2010). This represents a significant advancement in early-stage disease detection and large-scale monitoring, where timely intervention is crucial for mitigating potential crop yield loss.
@@
Despite the high classification accuracies achieved by the models, there were several limitations of the study. Positive class precision was consistently lower than negative class precision for models classifying RSD in Q200 and Q208, indicating a higher rate of false positives across the algorithms. This may have arisen from the assumption that RSD infection in a block was uniformly distributed by labeling all pixels within a block with the overall disease status of that block. This could result in regions with little or no RSD being mislabeled as ’RSD Positive’, thereby affecting model training. Future studies should avoid this assumption where possible and implement a higher resolution sampling method. Interestingly, this observation coincides with the only two varieties identified by Sugar Research Australia as partially resistant to RSD (Sugar Research Australia, 2021, 2024). Partial resistance might reduce the spread of RSD sufficiently to create patches of non-infected areas within otherwise infected blocks. However, to verify this hypothesis, further research with higher resolution sampling methods would be necessary.
This study focused on developing models and a prototype system for RSD detection within the Herbert region in February 2022. Future work should assess the influence of other environmental and agronomic predictors that could not be accounted for in this study to ensure broader applicability in the industry. This includes exploring the impacts of factors such as temperature, humidity, sunlight duration, flowering, other sugarcane varieties, simultaneous disease infections, and precipitation on the performance of classification models. Furthermore, RSD detection should be trialled at different growth stages of sugarcane using time-series data to improve its applicability to the industry. Given their proven success in previous time-series classification tasks, exploring the use of techniques such as recurrent neural networks or transformers for this application represents a promising direction for future research. Additionally, comparing various principal component analysis algorithms could further optimise vegetation index characteristics while addressing and mitigating issues such as overfitting.
@@ 5 Conclusion
In this study, we demonstrated that RSD can be classified for several varieties with ML algorithms from freely available satellite-based spectroscopy data. The best-performing machine learning algorithm was Support Vector Machine with a Radial Basis Kernel achieving between 85.64% and 96.55% accuracy depending on the variety. The performance of all machine learning algorithms was found to be significantly better than the null distribution for all varieties except SRA14. Inline with current literature, the inclusion of sugarcane variety information and vegetation indices improved classification rates of disease. Additionally, several varieties were found to be more challenging to classify for RSD compared to others. These promising initial results, coupled with the efficiency of classifying 76 blocks within minutes as opposed to months, indicates the potential benefits of implementing a large-scale health monitoring system using satellites and machine learning. Given that RSD is one of the most impactful diseases in sugarcane cultivation, this research significantly advances sugarcane disease management, with substantial practical applications. Providing farmers with a more accurate understanding of where RSD is present enables them to alter harvester paths to minimise disease spread, and source alternative clean, disease-free seeds from other blocks, or reliable suppliers, to ensure healthier crop cycles. Additionally, this will allow the implementation of strict hygiene protocols on infected blocks to prevent the disease from contaminating others. These applications indicate the potential of this research to support sustainable and efficient sugarcane cultivation.
@@ 7 Funding
Appendix D
Table 7: Summary of statistical significance from t-tests comparing healthy and diseased samples across different sugarcane varieties.’.’: adjusted p-value > 0.05, ’*’: 0.05 >= adjusted p-value > 0.01, ’**’: adjusted p-value <= 0.01
Vegetation Index Q200 Q208 Q240 Q253 SRA14
ARVI . . ** ** .
DWSI1 . * ** ** .
DWSI2 ** . . ** .
DWSI3 ** ** ** . **
DWSI4 . . ** ** .
DWSI5 . . ** ** .
NDMI . . ** ** .
NDVI . . ** ** .
NDWI . . ** ** .
NGRDI . . ** ** .
PSRI . ** ** . .
RVI . . ** ** .
SRI . . ** ** .
SR . . ** ** .
VARI . . ** ** .
Green Blue NDVI . . ** ** .
DWSI 6 ** ** ** . *
DWSI 7 ** ** ** . .
DWSI 8 . . ** ** .